OpenStudy (anonymous):

In a saturated solution of the salt MX, the concentrations of both M1+ and X1- are found to be 1*10^-3 M. What is Ksp for MX

OpenStudy (anonymous):

The Ksp expression for a salt is the product of the concentrations of the ions, with each concentration raised to a power equal to the coefficient of that ion in the balanced equation for the solubility equilibrium.

OpenStudy (anonymous):

so for M and X i can consider as X^2 and set it equal to 1*10^-3 M and solve for X?

OpenStudy (anonymous):

Yes that's right

OpenStudy (anonymous):

it keeps giving me a diff answer. the right answer is 1*10^-6

OpenStudy (sweetburger):

wait wait wait maybe i am just confused but wouldn't it be like this? Ksp= [1*10^-3][1*10^-3] this is usually how you find the Ksp given the molar solubility

OpenStudy (somy):

take a look at this pic and read carefully

OpenStudy (somy):

MX <----> M1+ X1 thus Ksp= [M1][X1] since concentrations for both M1 and X1 is same Ksp= [1*10^-3][1*10^-3] = if you are not allowed to use calculator then \[1\times1=1\] and now will work with powers, since its multiplication, powers are added up so \[-3+(-3)= -3-3=-6\] thus your answer will be \[1\times 10^-6\]

OpenStudy (anonymous):

ok thx

OpenStudy (somy):

:)